OpenStudy (anonymous):

Piecewise function help? Evaluate the piecewise function at the indicated values from the domain: | |x|, if x < -1 f(x) = { x^2, if -1 <(or equal to) x < 4 at x = 2 | -x, if x >(or equal to) 4

OpenStudy (anonymous):

(a.) f(-2) = 2 (b.) f(-2) = -4 (c.) f(-2) = -2 (d.) f(-2) = 4

OpenStudy (anonymous):

I'm pretty sure this has something to do with plugging in the numbers to see which statements are true but I'm not sure

OpenStudy (anonymous):

dont believe a is correct. 2 is bigger than -1

OpenStudy (anonymous):

why is c correct though

OpenStudy (anonymous):

I am having hard time figuring out the question, but should you just plug in a number and that will give you answer? -2^2=4

OpenStudy (anonymous):

you've claimed 3 answers so far and never provided any reasoning why makayla, please stop commenting on this

OpenStudy (anonymous):

Explain your reasoning then

OpenStudy (anonymous):

okay you wrote -2^2=4 but theres a - in it so -2^2=-4

OpenStudy (adamaero):

@JoeJoldin What function of the three do you use for x = -2 ?

OpenStudy (adamaero):

Which one fits the right "parameter"

OpenStudy (anonymous):

All you have to do is plug in x and if its true, then its correct.

OpenStudy (anonymous):

i'm down to b or c, still not sure which is which

OpenStudy (anonymous):

b

OpenStudy (anonymous):

does anybody else have an answer?

OpenStudy (anonymous):

No, -2^2=4 not -4

OpenStudy (anonymous):

so the answer is D? @TMA76

OpenStudy (adamaero):

@JoeJoldin What function of the three do you use for x = -2 ?

OpenStudy (anonymous):

The first??

OpenStudy (anonymous):

@adamaero

OpenStudy (adamaero):

yes, so f(-2) = ?

OpenStudy (adamaero):

but why do you have x = 2 when you posted the whole question?

OpenStudy (anonymous):

f(-2) = -4? I'm not sure how this is leading me to the answer

OpenStudy (anonymous):

@adamaero

OpenStudy (adamaero):

if x = -2, you said to use the first function which is the absolute value of x or |x| |-2| = 2 BUT in you original question you put x = 2 if it is x = 2, then you need to pick a different function Do you understand what I mean by "function?"

OpenStudy (anonymous):

No, I'm not understanding this, do you know what the answer is and can explain it? @adamaero

OpenStudy (adamaero):

forget about your specific problem for a second and take a look see: http://www.webgraphing.com/images/examples_piecewise_compound/ex2.gif

OpenStudy (adamaero):

see it?

OpenStudy (adamaero):

also, it may have been confusing when Makayla5646 spoke -2^2 = 4 (so it DOES equal 4) squared numbers loose their negative sign: -2*-2= (-1)(-1)(2)(2) = (1)(4) = 4 although -2^3 = -8 -x^(odd number) = -x*x*x...etc -x^(even number) = x*x*x*x...etc (negative sign forget)

OpenStudy (anonymous):

so the answer is D: f(-2) = 4 ?

OpenStudy (anonymous):

@adamaero

OpenStudy (adamaero):

nope, did you see the unrelated picture I posted?

OpenStudy (anonymous):

Yes and it didn't help me understand the answer

OpenStudy (adamaero):

you didn't understand the graph with piecewise?

OpenStudy (anonymous):

No. If D's not the answer and A doesn't work is it B or C? @adamaero

OpenStudy (adamaero):

Ok, forget about your specific problem. Once you understand how piecewise graphs work everything else will click I will break this unrelated picture down into three parts: http://www.webgraphing.com/images/examples_piecewise_compound/ex2.gif

OpenStudy (adamaero):

the first line, farthest to the left is the line of x + 3 (but only a part of it~ a piece)

OpenStudy (anonymous):

Forget it, closing the question and asking again since nobody is helping.

OpenStudy (adamaero):

ok. I don't give answers. I will guide you to the correct solution or check your answer. Only your effort to want to learn what's really going on will make the difference in understanding the concept here